OpenStudy (anonymous):

PLEASE HELP! WILL GIVE MEDAL! what is the simplified form of each expression? n^6/n^2 A)1/n^4 B)n^4 C)n^8 D)n^12

OpenStudy (***[isuru]***):

hi, use this law of indices... \[\frac{ n^{a} }{ n^{b}} = n^{a - b}\] what will be the answer ? :-)

OpenStudy (anonymous):

no idea. @raechelvictoria could u help me please!

OpenStudy (anonymous):

n^6 / n^2 When dividing numbers with the same base but different exponents, we subtract the exponent of the numerator from the exponent of the denominator, and raise the number to that result. What would the answer be then?

OpenStudy (anonymous):

@raechelvictoria I got c but I don't know if its right

OpenStudy (anonymous):

can you check it for me please

OpenStudy (anonymous):

No its B actually, because 6-2 = 4. Therefore n^4

OpenStudy (anonymous):

can you PLEASE help me with this one! (3x/2)^4 A)81x^4 B)6x^4 C)12x^4/8 D)81x^4/2 @raechelvictoria im really having trouble with this one!

OpenStudy (anonymous):

Distribute the exponent to all terms. (3x/2)^4 = (3^4)(x^4) / 2^4 Can you get the answer?

OpenStudy (anonymous):

@raechelvictoria would it be d?

OpenStudy (anonymous):

Its supposed to be 81x^4 / 16....

OpenStudy (anonymous):

@raechelvictoria its a. there is a divide by 16 on it I just accidentally didn't put it. sorry! can you check this form me? c^8d^-12/c^-4d^-8 A)c^12 d^4 B)c^-4 d^-4 C)d^-4/d^-12 D)c^12/ d^4 I got b